OpenStudy (anonymous):

Solve for y. 5x – y = 3 so is this were u bring 5x or to the other side of 3?

OpenStudy (anonymous):

-y=-5x+3 y=5x-3

jimthompson5910 (jim_thompson5910):

5x – y = 3 -y = 3 - 5x y = (3 - 5x)/(-1) y = 3/(-1) - (5x)/(-1) y = -3 + 5x y = 5x - 3

OpenStudy (anonymous):

Yes just move 5x to the other side and multiply by -1